OpenStudy (anonymous):

given f(x)=10-2x,find f(7) A. –4 B. 3 C. 7 D. 56

OpenStudy (mrnood):

this is the same as y=10-2x put x=7 into that equation and find the value for y

OpenStudy (anonymous):

replace x with 7; so f(7) = 10 - 2 *7 f(7)=10-14=-4 It's a).

OpenStudy (anonymous):

thanks